# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@
Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions
Martin G. Herold Max Planck Institute for Informatics, Saarland Informatics Campus    Danupon Nanongkai    Joachim Spoerhase University of Liverpool    Nithin Varma University of Cologne    Zihang Wu
@@
An important open problem is whether a $(1+\epsilon)$-approximation data structure exists. This is not known for any norm, even with higher (e.g. $\mathrm{poly}(n)\cdot o(d)$) space and query time. In this paper, we answer this question affirmatively. We present $(1+\epsilon)$-approximation data structures with the following guarantees.
@@ 1 Introduction
In the so-called low-dimensional regime [ANN+17]—when $d=o(\log n)$ and thus the data structure’s space and time complexities can be exponential in $d$—classical data structures can solve this problem efficiently (e.g. [Cla99, KR02, KL04, BKL06]). Recently, novel techniques such as locality-sensitive hashing and sketching (e.g. [IM98, AI08, AINR14, AR15, IW18]) have led to a new wave of data structures in the high-dimensional regime that solve the problem approximately: for any $\alpha\geq 1$, a $\alpha$-approximation data structure is the one that returns a point $c\in C$ whose distance to $q$ is at most $\alpha$ times the distance between $c^{*}$ and $q$. These recent data structures typically require space and time complexities that are linear or super linear in $d$, which is acceptable in most applications. In fact, many papers start with an assumption that $d=O(\log n)$ which can be assumed due to Johnson-Lindenstrauss lemma [JL84] at the cost of additional $\tilde{O}(d)$ space and query time.
@@
Scenarios like the above are a motivation behind at least two lines of research in machine learning. (i) In Feature Selection, the goal is to select a subset of relevant features of the input dataset. For example, Boutsidis, Drineas, and Mahoney in NeurIPS’09 [BDM09] (also see [BZMD15]) show that given a set $S\subseteq\mathbb{R}^{d}$ and a parameter $k\ll d$, we can select roughly $k$ coordinates such that we are required to only access these coordinates when computing $k$-means clustering, at the expense of a $(3+\epsilon)$ multiplicative approximation factor. The approximation factor was subsequently improved to $(1+\epsilon)$ in the STOC’15 paper by Cohen et al. [CEM+15]. (ii) In Explainable Clustering, given data that are already classified into $k$ clusters via cluster centers, the goal is to construct a decision tree with some desired properties that can classify the input data points into clusters without too much loss in the clustering cost compared to the original clustering. For example, Makarychev and Shan [MS21] show that given a $k$-median clustering on $\ell_{1}$ distance, one can construct such a decision tree that returns a clustering whose cost is $\tilde{O}(\log k)$ times the original cost. Since the number of accessed coordinates intuitively plays an important role in explainability, it is not surprising that one of the desired properties of the decision tree is that it never accesses more than $O(k)$ coordinates.
@@
  • For $\ell_{1}$-norm distances, an $O(\log(n))$-approximation data structure with $O(n\log(n)+n\log d)$ space and $O(n)$ query time can be obtained by adapting the explainable clustering algorithms of [GPSY23, MS23]. This is because these algorithms only require the center of each cluster, and not the remaining input data points. Weaker results can also be obtained from [EMN22, MS21, MDRF20, GJPS21].
  • For $\ell_{2}$-norm distances, a $(\sqrt{3}+\epsilon)$-approximation data structure with $\tilde{O}(n\log(d)/\epsilon^{2})$ space and $O(n\log(n)/\epsilon^{2})$ query time can be obtained by reducing to feature selection algorithms [BDM09, BZMD15, CEM+15] (see Appendix C). Weaker results can also be obtained from explainable clustering algorithms of [MS22, MS21, EMN22, GJPS21, CH22, MDRF20]
  • For $\ell_{p}$-norm distances, an $O(n^{p-1}\log^{2}(n))$-approximation data structure with $O(n\log(n)+n\log(d))$ space and $O(n)$ query time can be obtained from the explainable clustering algorithms of [GJPS21].
@@
We first observe that every $(1+\epsilon)$-approximate nearest neighbor data structure requires $\Omega(n)$ space and must make $\Omega(n)$ queries (see Theorem D.1). Our main contributions are data structures with nearly matching space and query bounds in the ultra-high dimensional regime. Our data structures require polynomial preprocessing time, answer each query correctly with $1-\delta$ probability, and guarantee the following bounds.
@@
  • For $\ell_{2}$-norm distances: $O(n\log^{2}(n/\delta)\log(d)/(\epsilon^{6}\delta^{4}))$ space and $O(n\log(n/\delta)/(\epsilon^{3}\delta^{2}))$ query time. See Theorem 5.17.
  • For $\ell_{p}$-norm distances: $O(n^{2}\log(n/\delta)\log(d)(\log\log n)^{p}/\epsilon^{p+2})$ space and $O(n\log(n/\delta)(\log\log n)^{p}/\epsilon^{p+2})$ query time. See Theorem 6.1.
@@
A recent line of research investigates explainable clustering [EMN22, MS21, MDRF20, GJPS21, GPSY23, MS23]. The input is a set $C$ of $k$ cluster centers. The output is a (random) decision tree assigning any query point to a cluster represented by a label in $[k]$. The current best result [GPSY23] for explainable $k$-median clustering under $\ell_{1}$-distances guarantees that for any set $P$ of data points the clustering generated by the decision tree is in expectation within a factor $O(\log k)$ of the optimal assignment of $P$ to the centers in $C$, that is, of the nearest-center assignment. The decision tree has $k$ leaves. Therefore, it can be interpreted as $O(\log k)$-approximate representation of the clustering $C$ with sublinear $\tilde{O}(k)$ space and query complexity. There are two aspects why this representation is considered explainable: (i) the decision tree is easy to interpret for humans (ii) it is compact as it has only sublinear $\tilde{O}(k)$ space and query complexity. When applied to a single query point these data structures constitute a $O(\log k)$-approximate nearest neighbor data structure in expectation.
Our techniques have implications in the converse direction. We argue (see Section 7) that our techniques give a sublinear $\tilde{O}(k)$ space and query complexity data structure for an $O(1)$-approximate nearest neighbor in expectation. This implies a sublinear $O(1)$-approximate representation for $k$-median clustering improving upon the $O(\log k)$-approximation obtained via decision trees. Notice that while our data structure does not produce decision trees it does constitute a compact representation of the clustering. Our techniques also imply a sublinear $O(1)$-approximate representation of $k$-means clustering in Euclidean space improving upon the $O(\log^{2}k)$-bound implied by explainable $k$-means clustering [MS22].
@@
As another example of where our techniques are applicable, consider the orthogonal range search problem. Here, we are given a set $C\subseteq\mathbb{R}^{d}$ to preprocess and need to answer the following query: Given $q_{1}$ and $q_{2}$ in $\mathbb{R}^{d}$, return all points $p\in C$ such that the $i^{th}$ attribute of $p$ has value between the $i^{th}$ attributes of $q_{1}$ and $q_{2}$, for every $i$. This problem can be solved using range trees and $k$-$d$ trees [Ben80, McC85, Cha86, Cha88] in the low-dimensional regime. For higher dimensions, some approximation algorithms have been proposed where some other points can be returned as long as they are not “far” from the query [AM00, CLM08, dFM10], i.e., the distance between the point returned and the range is within an $\epsilon$ fraction of the distance between $q_{1}$ and $q_{2}$—we call such a range search as $(1+\epsilon)$-approximate range search. We approach this problem, as before, from a data-structure point of view. Using similar techniques, we can design a data structure of size $\tilde{O}(n)$ that processes $n$ centers such that given any ‘valid’ orthogonal range $(q_{1},q_{2})$ (i.e., a range which contains at least one center point), queries only $\tilde{O}(n/\epsilon^{2})$ coordinates of $q_{1}$ and $q_{2}$ to report all centers that are within the $(1+\epsilon)$-approximate range. The details of this can be found in Appendix B.
@@
Section 2 gives an overview of the main technical contributions of the paper. Section 4 contains the outline of a simple data structure with space $\tilde{O}(n^{2})$ and query time $\tilde{O}(n)$ for $\ell_{1}$-norm, with some simplifying assumptions as warmup. Section 5 describes our $(1+\epsilon)$-approximate nearest neighbor data structures with $\tilde{O}(n)$ space and query complexity for the $\ell_{1}$- and $\ell_{2}$-norm. Section 6 describes our $(1+\epsilon)$-approximate nearest neighbor data structure with space $\tilde{O}(n^{2})$ and query time $\tilde{O}(n)$ that works for general $\ell_{p}$-norms. Finally, the implications of our results for other geometric data structure problems and some lower bound results are presented in the appendices.
@@ 2 Technical Overview
Let $c$ be an input point, let $q$ be an unknown query point, and let $\|c-q\|_{1}=\sum_{b}|c^{(b)}-q^{(b)}|$ denote their $\ell_{1}$ distance. First, observe that approximating this distance requires memorizing all coordinates of $c$ because $c$ and $q$ may differ in only a single coordinate. Our strategy is to compare distances to multiple input points without knowing the (approximate) distances. For intuition, consider when we have two input points $c_{1}$ and $c_{2}$. Consider an extreme case when $c_{1}$ and $c_{2}$ differ only in one coordinate $b$. In this case, keeping only this coordinate suffices to tell for any query $q$ if it is closer to $c_{1}$ or $c_{2}$. To extend this intuition to when $c_{1}$ and $c_{2}$ differ in many coordinates, a natural idea is to sample each coordinate $b$ independently with probability proportional to the difference $|c_{1}^{(b)}-c_{2}^{(b)}|$; more precisely, the probability that we sample coordinate $b$ is [^2] This can be viewed as an application of so-called importance sampling—a technique for variance reduction. The independent sampling process described here differs slightly from the one in Section 4.
@@
Under what we call bounding box assumption (which we will remove later), where $\min(c_{1}^{(b)},c_{2}^{(b)})\leq q^{(b)}\leq\max(c_{1}^{(b)},c_{2}^{(b)})$ for each coordinate $b$, we can find a $(1+\epsilon)$-approximate nearest neighbor between $c_{1}$ and $c_{2}$. This is due to the following property. (Details in Section 4.)
@@
With this idea, we can construct an $\tilde{O}(n^{2})$-space data structure by repeating the above for every pair $(c_{i},c_{j})$ of input points. It is not hard to find the approximately nearest neighbor to $q$ with $n-1$ comparisons: this is very simple if the comparator is exact, and the same idea can be extended for our approximate comparator as well; see Section 4 for details (also see [AFHN16] for previous work on imprecise comparisons).
@@
A more unclear point is: (ii) How many coordinates do we sample in total? For example, why is it not possible that in the worst case every pair $c_{i}$ and $c_{j}$ needs different coordinates and thus we are forced to select $\Omega(n^{2})$ coordinates in total. We show that this bad situation will not happen. To give an idea, we consider the special case $C\subseteq\{0,1\}^{d}$. The overall expected number of coordinates sampled is $S(C)=\sum_{b}p^{(b)}=\sum_{b}\max_{ij}p_{ij}^{(b)}$. (We ignore sampling repetition, which increases our bound only by $O(\log n)$.) Assume w.l.o.g. that $(c_{1},c_{2})$ is the closest pair in $C$. Notice that $p^{(b)}=p_{12}^{(b)}$ for every coordinate $b$ in the set $D$ of coordinates where $c_{1}$ and $c_{2}$ differ; they minimize $\|c_{i}-c_{j}\|_{1}$ and hence maximize $p_{ij}^{(b)}$. Therefore $S(C)$ drops by at most $\sum_{b}p_{12}^{(b)}=\sum_{b}|c_{1}^{(b)}-c_{2}^{(b)}|/\|c_{1}-c_{2}\|_{1}=1$ if we remove the coordinates in $D$ from all input points. On the other hand, we can identify $c_{1}$ and $c_{2}$ after removing $D$ obtaining a set $C^{\prime}$ with at most $n-1$ points. Thus $S(C)\leq S(C^{\prime})+1$. Using induction we can upper bound $S(C)$ by $n$. This inductive proof strategy can be extended to the general $\ell_{1}$ metric although the proof is notably more technical. Even though $\ell_{1}$ and $\ell_{2}$ metrics are equivalent on the Boolean cube, we do not know how to extend the above proof strategy to the general $\ell_{2}$ metric. To obtain a linear bound for $\ell_{2}$ we rely on an entirely different, linear-algebraic approach leveraging a latent connection between our sampling probabilities and the singular-value decomposition. Our proof is based on a connection to the feature selection approach by Boustidis et al. [BZMD15] for $k$-Means clustering. In fact, our sampling probabilities can be upper bounded by theirs, which are based on SVD. Notice, however, that we achieve a $(1+\epsilon)$-approximation for nearest neighbor as compared to their $(3+\epsilon)$-approximation for $k$-Means [^3] In Section C, we provide a formal reduction to feature selection by Boustidis et al. [BZMD15]. There is follow-up work achieving even guarantee $1+\epsilon$ for feature selection in clustering [CEM+15]. However, we do not know how to apply their approach to our setting. In particular, the above reduction is not applicable..
@@
Let $I$ denote the set of coordinates sampled by our global scheme. Memorizing these coordinates for all input points would still require $n|I|=\tilde{O}(n^{2})$ space. To obtain $\tilde{O}(n)$ space we apply standard dimension reduction tools such as the Johnson-Lindenstrauss transform for $\ell_{2}$ or the technique by Indyk [Ind06] for $\ell_{1}$. Specifically, we sample a suitable random $m\times|I|$ matrix $M$, where $m=O(\log n)$, and store the set $C^{\prime}=\{\,MSc_{i}^{(I)}\mid c_{i}\in C\,\}$ requiring $mn=\tilde{O}(n)$ space. Storing $M$ itself requires $m|I|=\tilde{O}(n)$ space, too, as does storing $I$ itself. Upon receiving a query $q$, we compute $MSq^{(I)}$ and output the nearest neighbor in $C^{\prime}$, which takes $\tilde{O}(n)$ coordinate queries. Here we use that the comparator property guarantees $\|Sq^{(I)}-Sc_{j}^{(I)}\|_{1}>(1+\epsilon/2)\|Sq^{(I)}-Sc_{i}^{(I)}\|_{1}$ if $\|q-c_{j}\|_{1}>(1+\epsilon)\|q-c_{i}\|_{1}$. Therefore, if we bound the distance distortion of the dimension reduction by, say, $101+\epsilon/$ 10 then $F(MSq^{(I)},MSc_{j}^{(I)})>(1+\epsilon/5)F(MSq^{(I)},MSc_{i}^{(I)})$ for the function $F$ from the $\ell_{1}$ dimension reduction by Indyk [Ind06]. This is sufficient for approximately comparing the distances.
@@
A natural attempt to resolve this issue is to “truncate” the query point so it comes sufficiently close to the bounding box. This approach is described in Section 6 and can in fact be applied to general $\ell_{p}$ metrics. However, it requires explicitly storing the projected points and therefore $\tilde{O}(n^{2})$ space.

### v2

@@
Abstract
@@
An important open problem is whether a $(1+\epsilon)$-approximation data structure exists. This is not known for any norm, even with higher (e.g. $\mathrm{poly}(n)\cdot o(d)$) space and query time.
In this paper, we answer this question affirmatively. We present $(1+\epsilon)$-approximation data structures with the following guarantees.
@@
keywords:
sublinear data structure, approximate nearest neighbor
@@ 1 Introduction
In the so-called low-dimensional regime [5]—when $d=o(\log n)$ and thus the data structure’s space and time complexities can be exponential in $d$—classical data structures can solve this problem efficiently (e.g. [16, 27, 28, 9]). Recently, novel techniques such as locality-sensitive hashing and sketching (e.g. [24, 2, 3, 6, 25]) have led to a new wave of data structures in the high-dimensional regime that solve the problem approximately: for any $\alpha\geq 1$, a $\alpha$-approximation data structure is the one that returns a point $c\in C$ whose distance to $q$ is at most $\alpha$ times the distance between $c^{*}$ and $q$. These recent data structures typically require space and time complexities that are linear or super linear in $d$, which is acceptable in most applications. In fact, many papers start with an assumption that $d=O(\log n)$ which can be assumed due to Johnson-Lindenstrauss lemma [26] at the cost of additional $\tilde{O}(d)$ space and query time.
@@
Scenarios like the above are a motivation behind at least two lines of research in machine learning. (i) In Feature Selection, the goal is to select a subset of relevant features of the input dataset. For example, Boutsidis, Drineas, and Mahoney in NeurIPS’09 [10] (also see [11]) show that given a set $S\subseteq\mathbb{R}^{d}$ and a parameter $k\ll d$, we can select roughly $k$ coordinates such that we are required to only access these coordinates when computing $k$-means clustering, at the expense of a $(3+\epsilon)$ multiplicative approximation factor. The approximation factor was subsequently improved to $(1+\epsilon)$ in the STOC’15 paper by Cohen et al. [17]. (ii) In Explainable Clustering, given data that are already classified into $k$ clusters via cluster centers, the goal is to construct a decision tree with some desired properties that can classify the input data points into clusters without too much loss in the clustering cost compared to the original clustering. For example, Makarychev and Shan [30] show that given a $k$-median clustering on $\ell_{1}$ distance, one can construct such a decision tree that returns a clustering whose cost is $\tilde{O}(\log k)$ times the original cost. Since the number of accessed coordinates intuitively plays an important role in explainability, it is not surprising that one of the desired properties of the decision tree is that it never accesses more than $O(k)$ coordinates.
@@
  • For $\ell_{1}$-norm distances, an $O(\log(n))$-approximation data structure with $O(n\log(n)+n\log d)$ space and $O(n)$ query time can be obtained by adapting the explainable clustering algorithms of [22, 32]. This is because these algorithms only require the center of each cluster, and not the remaining input data points. Weaker results can also be obtained from [19, 30, 34, 20].
  • For $\ell_{2}$-norm distances, a $(\sqrt{3}+\epsilon)$-approximation data structure with $\tilde{O}(n\log(d)/\epsilon^{2})$ space and $O(n\log(n)/\epsilon^{2})$ query time can be obtained by reducing to feature selection algorithms [10, 11, 17] (see Appendix F). [^2] In Appendix F we also argue why a natural reduction to feature selection [17] fails to give a $(1+\epsilon)$-approximate nearest neighbor data structure. Weaker results for the case of $\ell_{2}$-norm distances can also be obtained from explainable clustering algorithms of [31, 30, 19, 20, 12, 34].
  • For $\ell_{p}$-norm distances, an $O(n^{p-1}\log^{2}(n))$-approximation data structure with $O(n\log(n)+n\log(d))$ space and $O(n)$ query time can be obtained from the explainable clustering algorithms of [20].
@@
We first observe that every $(1+\epsilon)$-approximate nearest neighbor data structure requires $\Omega(n)$ space and must make $\Omega(n)$ queries (see Theorem G.1). Our main contributions are data structures with nearly matching space and query bounds in the ultra-high dimensional regime. Our data structures require polynomial preprocessing time, answer each query correctly with $1-\delta$ probability, and guarantee the following bounds.
@@
  • For $\ell_{2}$-norm distances: $O(n\log^{2}(n/\delta)\log(d)/(\epsilon^{6}\delta^{4}))$ space and $O(n\log(n/\delta)/(\epsilon^{3}\delta^{2}))$ query time. See Theorem B.7.
  • For $\ell_{p}$-norm distances: $O(n^{2}\log(n/\delta)\log(d)(\log\log n)^{p}/\epsilon^{p+2})$ space and $O(n\log(n/\delta)\allowbreak(\log\log n)^{p}/\epsilon^{p+2})$ query time. See Theorem D.1.
@@
A recent line of research investigates explainable clustering [19, 30, 34, 20, 22, 32]. The input is a set $C$ of $k$ cluster centers. The output is a (random) decision tree assigning any query point to a cluster represented by a label in $[k]$. The current best result [22] for explainable $k$-median clustering under $\ell_{1}$-distances guarantees that for any set $P$ of data points the clustering generated by the decision tree is in expectation within a factor $O(\log k)$ of the optimal assignment of $P$ to the centers in $C$, that is, of the nearest-center assignment. The decision tree has $k$ leaves. Therefore, it can be interpreted as $O(\log k)$-approximate representation of the clustering $C$ with sublinear $\tilde{O}(k)$ space and query complexity. There are two aspects why this representation is considered explainable: (i) the decision tree is easy to interpret for humans (ii) it is compact as it has only sublinear $\tilde{O}(k)$ space and query complexity. When applied to a single query point these data structures constitute a $O(\log k)$-approximate nearest neighbor data structure in expectation.
Our techniques have implications in the converse direction. We argue (see Section C) that our techniques give a sublinear $\tilde{O}(k)$ space and query complexity data structure for an $O(1)$-approximate nearest neighbor in expectation. This implies a sublinear $O(1)$-approximate representation for $k$-median clustering improving upon the $O(\log k)$-approximation obtained via decision trees. Notice that while our data structure does not produce decision trees it does constitute a compact representation of the clustering. Our techniques also imply a sublinear $O(1)$-approximate representation of $k$-means clustering in Euclidean space improving upon the $O(\log^{2}k)$-bound implied by explainable $k$-means clustering [31].
@@
As another example of where our techniques are applicable, consider the orthogonal range search problem. Here, we are given a set $C\subseteq\mathbb{R}^{d}$ to preprocess and need to answer the following query: Given $q_{1}$ and $q_{2}$ in $\mathbb{R}^{d}$, return all points $p\in C$ such that the $i^{th}$ attribute of $p$ has value between the $i^{th}$ attributes of $q_{1}$ and $q_{2}$, for every $i$. This problem can be solved using range trees and $k$-$d$ trees [8, 33, 13, 14] in the low-dimensional regime. For higher dimensions, some approximation algorithms have been proposed where some other points can be returned as long as they are not “far” from the query [7, 15, 18], i.e., the distance between the point returned and the range is within an $\epsilon$ fraction of the distance between $q_{1}$ and $q_{2}$—we call such a range search as $(1+\epsilon)$-approximate range search. We approach this problem, as before, from a data-structure point of view. Using similar techniques, we can design a data structure of size $\tilde{O}(n)$ that processes $n$ centers such that given any ‘valid’ orthogonal range $(q_{1},q_{2})$ (i.e., a range which contains at least one center point), queries only $\tilde{O}(n/\epsilon^{2})$ coordinates of $q_{1}$ and $q_{2}$ to report all centers that are within the $(1+\epsilon)$-approximate range. The details of this can be found in Appendix E.
@@
Section 2 gives an overview of the main technical contributions of the paper. Section 4 describes our $(1+\epsilon)$-approximate nearest neighbor data structures with $\tilde{O}(n)$ space and query complexity for the $\ell_{1}$- and $\ell_{2}$-norm.
Appendix A contains the outline of a simple data structure with space $\tilde{O}(n^{2})$ and query time $\tilde{O}(n)$ for $\ell_{1}$-norm, with some simplifying assumptions as warmup. Appendix B contains missing proofs of Section 4 and extension to $\ell_{2}$ case. Appendix C describes our expected constant-approximate nearest neighbor data structure, and shows the connection to explainable clustering. Appendix D describes our $(1+\epsilon)$-approximate nearest neighbor data structure with space $\tilde{O}(n^{2})$ and query time $\tilde{O}(n)$ that works for general $\ell_{p}$-norms. Appendix E shows the implications of our results for Range Search problem. Appendix F shows the connection between approximate nearest neighbor and feature selection for clustering. Appendix G shows $\Omega(n)$ space and query tight lower bound.
@@ 2 Technical Overview
Let $c$ be an input point, let $q$ be an unknown query point, and let $\|c-q\|_{1}=\sum_{b}|c^{(b)}-q^{(b)}|$ denote their $\ell_{1}$ distance. First, observe that approximating this distance requires memorizing all coordinates of $c$ because $c$ and $q$ may differ in only a single coordinate. Our strategy is to compare distances to multiple input points without knowing the (approximate) distances. For intuition, consider when we have two input points $c_{1}$ and $c_{2}$. Consider an extreme case when $c_{1}$ and $c_{2}$ differ only in one coordinate $b$. In this case, keeping only this coordinate suffices to tell for any query $q$ if it is closer to $c_{1}$ or $c_{2}$. To extend this intuition to when $c_{1}$ and $c_{2}$ differ in many coordinates, a natural idea is to sample each coordinate $b$ independently with probability proportional to the difference $|c_{1}^{(b)}-c_{2}^{(b)}|$; more precisely, the probability that we sample coordinate $b$ is [^3] This can be viewed as an application of so-called importance sampling—a technique for variance reduction. The independent sampling process described here differs slightly from the one in Section A.
@@
Under what we call bounding box assumption (which we will remove later), where $\min(c_{1}^{(b)},c_{2}^{(b)})\leq q^{(b)}\leq\max(c_{1}^{(b)},c_{2}^{(b)})$ for each coordinate $b$, we can find a $(1+\epsilon)$-approximate nearest neighbor between $c_{1}$ and $c_{2}$. This is due to the following property. (Details in Section A.)
@@
With this idea, we can construct an $\tilde{O}(n^{2})$-space data structure by repeating the above for every pair $(c_{i},c_{j})$ of input points. It is not hard to find the approximately nearest neighbor to $q$ with $n-1$ comparisons: this is very simple if the comparator is exact, and the same idea can be extended for our approximate comparator as well; see Section A for details (also see [1] for previous work on imprecise comparisons).
@@
A more unclear point is: (ii) How many coordinates do we sample in total? For example, why is it not possible that in the worst case every pair $c_{i}$ and $c_{j}$ needs different coordinates and thus we are forced to select $\Omega(n^{2})$ coordinates in total. We show that this bad situation will not happen. To give an idea, we consider the special case $C\subseteq\{0,1\}^{d}$. The overall expected number of coordinates sampled is $S(C)=\sum_{b}p^{(b)}=\sum_{b}\max_{ij}p_{ij}^{(b)}$. (We ignore sampling repetition, which increases our bound only by $O(\log n)$.) Assume w.l.o.g. that $(c_{1},c_{2})$ is the closest pair in $C$. Notice that $p^{(b)}=p_{12}^{(b)}$ for every coordinate $b$ in the set $D$ of coordinates where $c_{1}$ and $c_{2}$ differ; they minimize $\|c_{i}-c_{j}\|_{1}$ and hence maximize $p_{ij}^{(b)}$. Therefore $S(C)$ drops by at most $\sum_{b}p_{12}^{(b)}=\sum_{b}|c_{1}^{(b)}-c_{2}^{(b)}|/\|c_{1}-c_{2}\|_{1}=1$ if we remove the coordinates in $D$ from all input points. On the other hand, we can identify $c_{1}$ and $c_{2}$ after removing $D$ obtaining a set $C^{\prime}$ with at most $n-1$ points. Thus $S(C)\leq S(C^{\prime})+1$. Using induction we can upper bound $S(C)$ by $n$. This inductive proof strategy can be extended to the general $\ell_{1}$ metric although the proof is notably more technical. Even though $\ell_{1}$ and $\ell_{2}$ metrics are equivalent on the Boolean cube, we do not know how to extend the above proof strategy to the general $\ell_{2}$ metric. To obtain a linear bound for $\ell_{2}$ we rely on an entirely different, linear-algebraic approach leveraging a latent connection between our sampling probabilities and the singular-value decomposition. Our proof is based on a connection to the feature selection approach by Boustidis et al. [11] for $k$-Means clustering. In fact, our sampling probabilities can be upper bounded by theirs, which are based on SVD. Notice, however, that we achieve a $(1+\epsilon)$-approximation for nearest neighbor as compared to their $(3+\epsilon)$-approximation for $k$-Means. [^4] In Section F, we provide a formal reduction to feature selection by Boustidis et al. [11]. There is follow-up work achieving even guarantee $1+\epsilon$ for feature selection in clustering [17]. However, we do not know how to apply their approach to our setting. In particular, the above reduction is not applicable.
@@
Let $I$ denote the set of coordinates sampled by our global scheme. Memorizing these coordinates for all input points would still require $n|I|=\tilde{O}(n^{2})$ space. To obtain $\tilde{O}(n)$ space we apply standard dimension reduction tools such as the Johnson-Lindenstrauss transform for $\ell_{2}$ or the technique by Indyk [23] for $\ell_{1}$. Specifically, we sample a suitable random $m\times|I|$ matrix $M$, where $m=O(\log n)$, and store the set $C^{\prime}=\{\,MSc_{i}^{(I)}\mid c_{i}\in C\,\}$ requiring $mn=\tilde{O}(n)$ space. Storing $M$ itself requires $m|I|=\tilde{O}(n)$ space, too, as does storing $I$ itself. Upon receiving a query $q$, we compute $MSq^{(I)}$ and output the nearest neighbor in $C^{\prime}$, which takes $\tilde{O}(n)$ coordinate queries. Here we use that the comparator property guarantees $\|Sq^{(I)}-Sc_{j}^{(I)}\|_{1}>(1+\epsilon/2)\|Sq^{(I)}-Sc_{i}^{(I)}\|_{1}$ if $\|q-c_{j}\|_{1}>(1+\epsilon)\|q-c_{i}\|_{1}$. Therefore, if we bound the distance distortion of the dimension reduction by, say, $101+\epsilon/$ 10 then $F(MSq^{(I)},MSc_{j}^{(I)})>(1+\epsilon/5)F(MSq^{(I)},MSc_{i}^{(I)})$ for the function $F$ from the $\ell_{1}$ dimension reduction by Indyk [23]. This is sufficient for approximately comparing the distances.
@@
A natural attempt to resolve this issue is to “truncate” the query point so it comes sufficiently close to the bounding box. This approach is described in Section D and can in fact be applied to general $\ell_{p}$ metrics. However, it requires explicitly storing the projected points and therefore $\tilde{O}(n^{2})$ space.